OpenStudy (anonymous):

What is the vertex of the function x2 - 6x + 8 = 0?

OpenStudy (ribhu):

x = -b/2a = 6/2 = 3 and put x =3 in the quadratic expression, and u will get y = -1

OpenStudy (ribhu):

i hope @roseromeroo this is of help to u

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html This should help you. :)

OpenStudy (anonymous):

Any future things.

OpenStudy (mikeymaximum):

*Multiply the 1st and last coefficient (1*8 = 8). Find two numbers that give you 8 when multiplied and -6 (middle coefficien) when added/subtracted. The numbers are: -2, -4 First: rewrite the equation with the new middle numbers > x^2 - 2x - 4x +8 = 0 Sec: group "like" terms and factor both sets of parenthesis > (x^2 - 2x) - (4x +8) = 0 x(x - 2) - 4(x - 2) = 0 (x - 2)(x - 4) = 0 Third: solve for both "x" variables by setting the parenthesis to equal "0" > x - 2 = 0 x - 2+2 = 0+2 x = 2 x - 4 = 0 x - 4+4 = 0+4 x = 4 (2, 4) Source(s):